OpenStudy (bunnykitten02):

MEDALS!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! PLEASE HELP What property is shown? –12 + (27 + 12) = –12 + (12 + 27) A. associative property B. commutative property C. distributive property D. opposite of a sum property

OpenStudy (bunnykitten02):

PLEASE HELP

OpenStudy (anonymous):

haha! this is the commutative property. All it takes to figure out is a simple search of the definitions of each property. ex: The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2.

OpenStudy (bunnykitten02):

thank you